# Continuous-time quantum walk-based ansätze on neutral atom hardware

### v1

@@ V Results: Coherent quenches of bracelet states
have identical populations ($|\langle z|\rho_{\rm inc}|z\rangle|^{2}=|\langle z|\rho_{\rm coh}|z\rangle|^{2}$), even though there are nontrivial off-diagonal matrix elements in $\rho_{\rm inc}$ and (if prepared exactly) none in $\rho_{\rm inc}$.
@@
Figure 10 illustrates that, at short evolution times, the dynamics of these two states are equivalent, but at longer times, interference effects cause the distributions to diverge. The coherent and incoherent perfect walk dynamics (gray) show the earliest divergence and clearly distinguish the coherent and incoherent states. Emulated neutral atom dynamics (black) follow the perfect walk dynamics, as expected from parameter matching the virtual and physical systems. However, at longer times the similarity diverges due to perturbative blockade violations and $R^{-6}$ Rydberg interactions outside of the blockade radius (see App. B). This suggests that, while for short times, such as state preparation, the two dynamics approximately match. For longer evolution times, the dynamic blockade radius approximation begins to break down, and the Rydberg atom Hamiltonian can no longer accurately represent CTQW on independent set subspaces.
@@
Figure 10: Quenches for target state $|z^{\rm half}\rangle=|[000101]\rangle$ under incoherent product state preparation (top) and bracelet state preparation (bottom), plotting the probability of measuring $|z^{\rm half}\rangle$ in the dihedrally symmetric subspace of $\mathcal{G}$. Aquila data averaged over $100$ shots per $\tau$.
@@ VI Conclusion
Unlike earlier approaches [2] that used quantum walks as variational routines for preparing unknown states, here we target nontrivial but efficiently describable entangled states [29] chosen as benchmarks for assessing interference and amplification on NISQ hardware. In particular, we focus on bracelet states $|[z]\rangle$ that respect the dihedral symmetry of an $N$-vertex ring, and product states $|z\rangle$ that must nonetheless have nontrivial dynamics due to the Hilbert space. Their entanglement does not arise from an explicit mixer term, but from the non-separable structure of the independent-set Hilbert space itself. We find that this constrained space, together with restricted local controls, is sufficient to support variational ansätze capable of preparing such structured target states.
We demonstrate a key use of quantum processors as analogue simulators of abstract models: the dynamics of CTQWs on independent-set graphs can be mapped directly onto the nonequilibrium dynamics of the Rydberg Hamiltonian. The independent-set constraint is naturally enforced by the Rydberg blockade, enabling a hardware representation of nonequilibrium dynamics on constrained subspaces that has previously only been explored in the context of quantum scar dynamics [22]. Despite substantial measurement errors, we find that Aquila can reproduce the same target states as the abstract CTQW system through direct parameter transfer with little to no further optimization. In the case of unentangled product-state targets, hardware circuits reproduce the hallmark amplification through constructive interference of efficient CTQW-inspired protocols. Scaling analysis shows super-linear growth in success probability with system size, and effective polynomial orders that are consistent with the super-quadratic convergence bounds expected from theory. Although this does not constitute evidence of an algorithmic speedup, the analysis demonstrates that the mechanisms underpinning such speedups are already observable on NISQ neutral atom hardware.
@@ Appendix C State Preparation Results
Tables 2 and 3 summarize the numerical and hardware results obtained for the preparation of product states and bracelet states, respectively. In both cases, the State column identifies the target configuration, where $N$ denotes the number of vertices in the constraint graph and $|\mathcal{V}|$ the number of vertices in the corresponding walk graph $\mathcal{G}$. The Depth ($p$) column records the number of ansatz layers. For product state preparation (Table 2), the variationally optimised parameters are the fiducial and walk times $\tau_{0}$ and $\tau_{1}$, and $J_{\rm eff}$ is the effective coupling (see Eq. 28) computed from the walk times. For bracelet states (Table 3), the relevant parameters are the effective evolution time $\tau_{\text{eff}}=\tau p$ and the phase jumps $\gamma$.

### v2

@@ V Results: Coherent quenches of bracelet states
have identical populations ($|\langle z|\rho_{\rm inc}|z\rangle|^{2}=|\langle z|\rho_{\rm coh}|z\rangle|^{2}$), even though there are nontrivial off-diagonal matrix elements in $\rho_{\rm coh}$ and (if prepared exactly) none in $\rho_{\rm inc}$.
@@
Figure 10 illustrates that, at short evolution times, the dynamics of these two states are equivalent, but at longer times, interference effects cause the distributions to diverge. The coherent and incoherent perfect walk dynamics (gray dashed) show the earliest divergence and clearly distinguish the coherent and incoherent states. Emulated neutral atom dynamics (purple) follow the perfect walk dynamics, as expected from parameter matching the virtual and physical systems. However, at longer times the similarity diverges due to perturbative blockade violations and $R^{-6}$ Rydberg interactions outside of the blockade radius (see App. B). This suggests that, while for short times, such as state preparation, the two dynamics approximately match. For longer evolution times, the dynamic blockade radius approximation begins to break down, and the Rydberg atom Hamiltonian can no longer accurately represent CTQW on independent set subspaces.
@@
Figure 10: Quenches for target state $|z^{\rm half}\rangle=|[000101]\rangle$, plotting the probability of measuring $|z^{\rm half}\rangle$ in the dihedrally symmetric subspace of $\mathcal{G}$. Top shows incoherent product state preparation. Bottom shows bracelet state preparation. Aquila data averaged over $100$ shots per $\tau$.
@@ VI Conclusion
Unlike earlier approaches [2] that used quantum walks as variational routines for preparing unknown states, here we target nontrivial but efficiently describable states [34] chosen as benchmarks for assessing interference and amplification on NISQ hardware. In particular, we focus on bracelet states that respect the dihedral symmetry of an $N$-vertex ring, and product states whose preparation nonetheless requires nontrivial dynamics through the constrained Hilbert space. For the bracelet states, entanglement emerges from the non-separable structure of the independent-set Hilbert space itself. We find that this constrained space, together with restricted local controls, is sufficient to support variational ansätze capable of preparing such structured target states.
We demonstrate a key use of quantum processors as analogue simulators of abstract models: the dynamics of CTQWs on independent-set graphs can be mapped directly onto the nonequilibrium dynamics of the Rydberg Hamiltonian. The independent-set constraint is naturally enforced by the Rydberg blockade, enabling a hardware representation of nonequilibrium dynamics on constrained subspaces that has previously only been explored in the context of quantum scar dynamics [18]. Despite substantial measurement errors, we find that Aquila can reproduce the same target states as the abstract CTQW system through direct parameter transfer with little to no further optimization.
For product-state targets, hardware circuits reproduce the hallmark amplification through constructive interference of efficient CTQW-inspired protocols. The $z^{\rm MIS}$ states benefit from perfect state transfer on an embedded hypercube subgraph, with ideal CTQW dynamics predicting near-unity amplification even at a single ansatz iteration. The $z^{\rm half}$ states provide a more representative benchmark. Here, our results suggest a conservative theoretical upper-bound on the achievable polynomial-order speedup in query complexity of $n\approx 8$ at ansatz depth one, rising to $n\approx 15$ at depth two and $n>50$ at depth three. On Aquila, we observe $n\approx 4$ at depth one and $n\approx 8$ at depth two, demonstrating that the fast-forward dynamics characteristic of efficient CTQW protocols are already accessible at low circuit depth. This opens the possibility of realizing super-quadratic amplification in algorithmic applications on current hardware.
For bracelet-state targets, our results extend the non-adiabatic shortcut demonstrated for product-state MIS preparation via sweep-quench-sweep protocols [28] to the preparation of symmetric entangled states, successfully applying a CTQW-based framework that predicts an advantage over adiabatic protocols. While the influence of Van der Waals interactions in the Rydberg dynamics precluded the direct transfer of CTQW-derived parameters, we found that it was nevertheless possible to identify parameters that are jointly optimal for both the ideal CTQW and the Rydberg evolution, with resulting walk times retaining the $\tau_{\rm eff}\propto 1/\Delta_{\rm min}$ scaling predicted by the non-adiabatic protocol. To our knowledge, this is the first attempt to prepare such states on hardware.
@@ Appendix C State Preparation Results
This appendix collects detailed numerical results for state preparation experiments. Table 4 reports the power-law fit parameters ($c$, $\alpha$, and $R^{2}$) for the amplification scaling $A=c\,|\mathcal{V}|^{\alpha}$ shown in Fig. 6. Tables 2 and 3 provide per-instance success probabilities for product states and bracelet states, respectively.
In both Tables 2 and 3, the State column identifies the target configuration, where $N$ denotes the number of vertices in the constraint graph and $|\mathcal{V}|$ the number of vertices in the corresponding walk graph $\mathcal{G}$. The Depth ($p$) column records the number of ansatz layers. For product state preparation (Table 2), the variationally optimised parameters are the fiducial and walk times $\tau_{0}$ and $\tau_{1}$, and $J_{\rm eff}$ is the effective coupling (see Eq. 28) computed from the walk times. For bracelet states (Table 3), the relevant parameters are the effective evolution time $\tau_{\text{eff}}=\tau p$ and the phase jumps $\gamma$.
@@
Target Platform $p$ $c$ $\alpha$ $R^{2}$
$z^{\rm half}$ CTQW 1 1.099 $0.878\pm 0.006$ 0.9870
2 1.112 $0.935\pm 0.003$ 0.9997
3 1.015 $0.981\pm 0.002$ 0.9997
Emulation 1 0.632 $0.877\pm 0.009$ 0.9881
2 0.719 $0.934\pm 0.006$ 0.9961
3 0.785 $0.930\pm 0.006$ 0.9959
Hardware 1 0.595 $0.729\pm 0.041$ 0.9326
2 0.463 $0.876\pm 0.025$ 0.9374
3 0.418 $0.734\pm 0.051$ 0.7889
$z^{\rm MIS}$ CTQW 1 0.995 $0.996\pm 0.001$ 0.9972
2 1.000 $1.000\pm 0.000$ 1.0000
Emulation 1 0.641 $1.000\pm 0.004$ 0.9838
2 0.507 $0.967\pm 0.006$ 0.9857
Hardware 1 0.512 $0.820\pm 0.029$ 0.9724
2 0.285 $0.703\pm 0.067$ 0.9250
Table 4: Power-law fit parameters for amplification scaling $A=c\,|\mathcal{V}|^{\alpha}$ (see Eq. 13) across CTQW (Perfect), noiseless Rydberg emulation, and Aquila hardware. Values are listed for depths $p=1,2,3$ for $z^{\rm half}$ and $p=1,2$ for $z^{\rm MIS}$).